OpenStudy (anonymous):

Find the values of x in the following equation: |9x + 9| = 18

OpenStudy (anonymous):

1 and -2

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

remove absolute values signs and write TWO equations

OpenStudy (anonymous):

\[9x+9=18\] \[9x+9=-18\] solve both of these to get your answer

OpenStudy (anonymous):

as stroogle said, first one gives \[x=1\] but second one give \[9x=-27\] \[x=-3\]

OpenStudy (anonymous):

your x is between -18 and +18 to get the absolute operator off the eq.

OpenStudy (anonymous):

Good way to deal with these is to square, then take the root of both sides of the equation. This gets rid of the absolute value, and gives: \[9x+9=\pm18\] giving 1 and -3 (I wrote the wrong answer above sorry)

OpenStudy (anonymous):

the root??? c'mon.

OpenStudy (anonymous):

\[-18\le9x+9\le18\]\[-27\le9x \le9\]\[-3\le x \le1\]

OpenStudy (anonymous):

look if \[|\text{blah}|=p\] then either \[\text{blah}=p\] or \[\text{blah}=-p\] that is all. remove absolute value, solve two equations. finito

OpenStudy (anonymous):

good lord another answer. @daomown question says EQUAL not less than. course if it said LESS THAN your answer would be right

OpenStudy (anonymous):

the point is that in order to solve the abs value you need to put those less thans so that way you are indicating your algebraic expression is between both integers (-18 and 18 in this case), of course the way it solves is that x = -3, and x=1

jhonyy9 (jhonyy9):

9x+9=18 and 9x+9=-18 9x=9 9x=-27 x=1 x= -27/9 = -3